OpenStudy (anonymous):

A rock takes 2.95 s to hit the ground when it is thrown straight up from the cliff with an initial velocity of 8.3 m/s t = 2.95 s v = 8.3 m/s (a):Calculate the height of the cliff in m. (b):How long would it take to reach the ground if it is thrown straight down with the same speed?

OpenStudy (anonymous):

fort the first one you can use the formula \(-dy=v_isin \theta- \frac{1}{2}gt^2\) i put "minus" because i set down as negative direction

OpenStudy (anonymous):

ok working that all out i got 18.16

OpenStudy (anonymous):

ok give me a sec, i'll try to solve it

OpenStudy (anonymous):

i'm not sure but for the second one, i got the same time O.o i also did another solution

OpenStudy (anonymous):

by same time you mean 2.95seconds

OpenStudy (anonymous):

when i put 2.95 seconds in it said it was wrong

OpenStudy (anonymous):

i told you, i'm not sure i'm still figuring out how to calculate it

OpenStudy (anonymous):

its no problem i appreciate all the help

OpenStudy (anonymous):

now i get it is it 1.25s?

OpenStudy (anonymous):

i still used the same formula, but since it is thrown downwards it's easier to set down as positive \(dy=vit+ \frac{1}{2}gt^2\) \(18.16=(8.3)t+ \frac{1}{2}(18.16)t^2\) \(0= 4.9t^2+(8.3)t-18.16\) where dy= 18.16 then use the quadratic formula to solve for t

OpenStudy (anonymous):

1.25 is also wrong with this formula do you have to take the squareroot?

OpenStudy (anonymous):

yeah, when you put in the quadratic formula, you need to take the square root of the discriminant sorry, i guess that's all i can think of T_T i'll ask someone to help you @wolfe8

OpenStudy (anonymous):

appreciate it